The melting point of pure gold is 1,945 Fahrenheit. Convert this temperature to Celsius using the formula, C=5/9(F - 32)
Please help and by the way 5/9 is a fraction Thanks

Answer by stanbon(75887)

The melting point of pure gold is 1,945 Fahrenheit. Convert this temperature to Celsius using the formula, C=(5/9)(F - 32)
----------
C(1945) = (5/9)(1945-32)
= (5/9)(1913)
= 1062.78 degrees Celsius
Cheers,
Stan H.